# Varco Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) necessitates robust evaluation methodologies. Current benchmarking approaches often rely on comparing model outputs against predefined prompts and reference outputs. Relying on predefined reference outputs hinders flexible adaptation of benchmarks to the rapidly evolving capabilities of LLMs. This limitation necessitates periodic efforts to prepare new benchmarks. To keep pace with rapidly evolving LLM capabilities, we propose a more flexible benchmarking approach. Our method, Varco Arena, provides reference-free benchmarking of LLMs in tournament style. Varco Arena directly compares LLM outputs across a diverse set of prompts, determining model rankings through a single-elimination tournament structure. This direct pairwise comparison offers two key advantages: (1) Direct comparison, unmediated by reference text, more effectively orders competing LLMs, resulting in more reliable rankings, and (2) reference-free approach to benchmarking adds flexibility in updating benchmark prompts by eliminating the need for quality references. Our empirical results, supported by simulation experiments, demonstrate that the Varco Arena tournament approach aligns better with the current Elo model for benchmarking LLMs. The alignment is measured in terms of Spearman correlation, showing improvement over current practice of benchmarking that use reference outputs as comparison anchors.
@@ 1 Introduction
The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks. The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios, exemplified by generalist assistants (Ouyang et al. 2022; Köpf et al. 2024; Roziere et al. 2023). As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant. In response to this challenge, numerous benchmarks evaluating LLM capability have emerged (Hendrycks et al. 2020; Srivastava et al. 2023; Zhong et al. 2024). Many LLM benchmarks employ formats amenable to automated scoring. Examples include benchmarks for arithmetic problems (Gao et al. 2022; Patel, Bhattamishra, and Goyal 2021), multiple-choice questions (Lin, Hilton, and Evans 2022), and code execution (Austin et al. 2021; Chen et al. 2021). While these benchmarks are valuable, they primarily assess problem-solving abilities. The need for LLM benchmark extends beyond this, as the primary value of LLMs lies in their generative capabilities. Researchers have primarily relied on pariwise comparison to evaluate and rank the generative capabilities of LLMs. A representative example of this approach is Chatbot Arena (Chiang et al. 2024), which computes Elo ratings based on a massive number of human-annotated pairwise comparisons. While Chatbot Arena benefited from the reliability of open-ended prompts from users, acquiring such a large number of annotations for benchmarking LLMs is time-consuming and resource-intensive. To address these challenges, several benchmark datasets for generation tasks have been developed. These leverage reference responses for ranking LLMs using automated LLM judges (Zheng et al. 2024) includes AlpacaEval (Li et al. 2023), Arena-Hard-Auto (Li et al. 2024), and MTBench (Zheng et al. 2023). Relying on reference outputs offers two key advantages: (1) The number of comparisons required for ranking multiple LLMs reduces linear to the number of LLMs. (2) Reference texts establish a quality bar for evaluating responses.
However, we propose that directly comparing the responses of LLMs, rather than relying on reference outputs, is a more effective way to achieve accurate LLM rankings. We suggest that combined with tournament-styled match making, direct comparison of outputs can achieve a better correlation to human preference within the same number of comparisons. Additionally, reference-free benchmarking mitigates potential biases in open-ended generation tasks and provides more flexibility compared to static benchmark datasets.
@@
Figure 1: Varco Arena directly compares LLM output pairs in single-elimination tournament structure rather than comparing LLM outputs to reference text. In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct comparison is more intuitive and results in better separability.
We devised Varco Arena, which directly compares LLM outputs without the need for reference text as a comparison baseline. We leveraged the efficiency of single-elimination tournaments, which always require a linear number of matches relative to the number of participants. For brevity, we will refer to single-elimination tournaments simply as tournaments throughout the paper. Varco Arena conducts a tournament for each prompt across all participating LLMs and then computes Elo ratings (Elo and Sloan 1978) based on the results of all matches occurred. By performing tournament matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs within a controlled number of matches. These win rates are then interpreted as Elo ratings, allowing us to create a comprehensive ranking of participating LLMs based on their relative performance.
We validate the effectiveness of our approach, Varco Arena, against the current practice of LLM benchmarks, anchored comparison to reference, through two experiments: one is unbiased judge simulation and the other is running Varco Arena with LLM-as-a-judge on Arena-Hard-Auto benchmark dataset. These experiments confirm that our reference-free benchmarking approach outperforms current practices in terms of rank reliability, given the same number of comparisons. Notably, the simulation experiment, which emulates outcomes under an unbiased judge and controls for factors such as the number of participant LLMs, test prompts, and judge precision, demonstrates that the empirical superiority of our reference-free approach is not coincidental. Instead, it reveals the effectiveness of direct comparison between LLM pairs compared to the current practice of mediated comparison with references for ranking LLMs.
To summarize, our key contributions are as follows:
@@
    We introduce Varco Arena, a reference-free benchmarking method for LLMs that employs direct comparison of outputs through a single-elimination tournament structure across test prompts.
@@
    We demonstrate, through conceptual validation and empirical results, that Varco Arena delivers more reliable rankings than current practices of comparing LLM outputs to reference texts, while requiring fewer comparisons.
@@ 2 Preliminaries: Quantifying Generation Ability
Quantifying an LLM’s generation ability is crucial for ranking models, but it presents several challenges. The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as test prompts and the subjective nature of human preferences. A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its capability in real-world scenarios. Two popular measures for LLM generation ability are: win rate against reference text and elo rating.
2.1 Win rate Against the Reference Text
Several benchmarks quantify response generation quality, with AlpacaEval (Li et al. 2023) and Arena-Hard (Li et al. 2024) being representative examples. These benchmarks employ LLMs as judges for automated evaluation of generation capability. Given a prompt and two responses, the judge LLM is tasked with determining the superior response. These benchmarks include predefined sets of test prompts and reference outputs that serve as quality standards. The evaluation process involves comparing the LLM’s output against the reference to assess response quality. The LLM’s win rate across test prompts is used as a measure of its generation proficiency.
@@ 2.2 Elo Rating
The Elo rating system, introduced by Elo and Sloan, has since become a popular method for quantifying performance levels in competitive sports and, more recently, for evaluating the generative abilities of LLMs. The primary purpose of the Elo rating system is to represent a participant’s skill level as a single scalar value, enabling the prediction of relative win rates between participants who have never directly competed. Assessing the superiority of one LLM over another in terms of generative capability shares several similarities with determining winners in competitive sports. Varying test prompts may yield different results, similar to how weather or other factors can influence the outcome of a sports match. While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently. With Elo ratings, one can expect the relative chance of winning between a1 pair of players. The expected win rate between a pair of players $i$ and $j$ is computed as follows:
@@ 3 Varco Arena
Varco Arena represents our novel approach to reference-free benchmarking of LLMs. Rather than comparing LLM responses to references, we propose iterated tournaments over benchmark dataset to build leaderboards. This addresses concerns about effectiveness and scalability of using direct comparison over traditional benchmarking practice of relying on references. Section 3.1 and Algorithm 1 detail how Varco Arena runs iterated tournaments as we suggest. We then explain our rationale for using direct comparison of LLM responses over bridged comparing to references (Section 3.2), and argue why iterated tournaments can produce effective LLM leaderboards (Section 3.3).
Algorithm 1 Running Iterated Tournaments over Benchmark for multiple LLMs
1:Test prompts $X$, LLMs $M$, outputs $O_{ij}$ = $m_{j}(x_{i})$
@@
4:     if $\text{IsBetter}(O_{x,m_{1}},O_{x,m_{2}})$ then
5:         return $m_{1}$
6:     else
7:         return $m_{2}$
8:     end $\triangleright$ omitted byes for conciseness
9:end function
10:function SingleElim(players, $x$, results)
11:     if ||||players|=2|=2| = 2 then
12:         $m_{\text{win}}$leftarrow Match(players[0], players[1], $x$)
13:         results.append($m_{\text{win}}$)
14:         return $m_{\text{win}}$
15:     end if
@@
17:     l_bracket $\leftarrow$ SingleElim(players[:mid], $x$, results)
18:     r_bracket $\leftarrow$ SingleElim(players[mid:], $x$, results)
19:     return SingleElim(l_bracket + r_bracket, $x$, results)
20:end function
21:function Tournament2Leaderboard($X,M$)
22:     results $\leftarrow$ []
23:     for $x_{i}\in X$ do
24:         SingleElim(Shuffled($M$), $x_{i}$, results)
25:     end for
26:     return ComputeElo(results)
27:end function
3.1 Iterating Tournaments over Benchmark Dataset to Elo ratings
Figure 1 and Algorithm 1 illustrate the methodology of Varco Arena for benchmarking Large Language Models (LLMs) using a tournament approach. In this context, $|X|$ denotes the cardinality of the prompt set in the benchmark dataset (i.e. number of prompts in the dataset). As described running Varco Arena invokes tournament of participant LLMs for each prompt in benchmark dataset, which we will refer to as iterated tournaments.
The application of tournament structures to LLM benchmarking presents both opportunities and challenges. While a single tournament effectively identifies a champion, it does not provide the granularity required for a comprehensive ranking system; it only declares a champion, leaving the precise ordering of other participants ambiguous. To retain the efficient nature of tournaments but to accomplish a fine-grained ranking of LLMs, what modifications should be applied to? To mitigate the aforementioned ranking resolution issue, we propose conducting independent tournaments for each benchmark prompt, with randomized initial match-ups. This approach, which we term iterated tournaments, offers several advantages:
@@ 3.1 Iterating Tournaments over Benchmark Dataset to Elo ratings
    It facilitates the resolution of ties among non-champion participants from previous tournaments.
@@
    It reduces the impact of unfavorable match-ups in any single tournament.
@@
    While not providing exhaustive pairwise comparisons, the accumulated matches between model pairs enable more precise estimation of relative win rates and Elo ratings, culminating in a well-aligned overall ranking.
In terms of total number of matches, our iterated tournaments approach requires $|X|*(n_{\text{models}}-1)$ which is always $|X|$ numbers of comparisons saved compared to anchored comparison to references. A favorable trait of conducting tournaments for each prompt is that every participating LLM engages in at least one competition per prompt. The number of matches an LLM faces per test prompt varies from 1 to $\lceil\text{log}_{2}(n_{\text{models}})\rceil$, based on its performance. Allocating more matches for better LLMs will help precise separation between higher rated models.
We further elaborate and argue why aggregating multiple tournaments should reach a good sorting of participant LLMs in Section 3.3. Additionally, we provide prospect match numbers LLMs would confront within, offering a comprehensive view of the method’s efficiency and effectiveness.
3.2 Comparing to Reference is not Always Helpful
Here, we inspect into failure modes of relying on references for ranking LLMs, which further elaborates the previously mentioned concerns on current practices for LLM benchmarking. Consider a highly accurate judge capable of definitively comparing two outputs of different quality without errors or ties. In this scenario, all matches involving participant LLMs contribute to their relative win rates and, ultimately, to their final Elo ratings (Section 3.3 and Equation 2). Consequently, all matches involving the LLM of interest contribute to Elo rating and ranking of the LLM.
When evaluating LLM performance relative to reference texts, even with the same precise judge, the comparison doesn’t always provide a reliable measure of relative win rates. Suppose we have only two participants of interest $\text{M}_{1}$ and $\text{M}_{2}$ and an instruction $X_{i}$ to prompt these LLMs. There might be two corresponding responses $\text{M}_{1}(X_{i})$, $\text{M}_{2}(X_{i})$ and also a reference text $Y_{i}$. Equation 2 illustrates four possible scenarios. In two of these cases, the match outcome is useful for comparing two LLMs, specifically when the reference text loses to one model and wins against the other, working as a useful anchor for comparison. In the other two cases, comparing to $Y_{i}$ fails to differentiate between $\text{M}_{1}(X_{i})$ and $\text{M}_{2}(X_{i})$.
$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ ~{}~{}~{}~{}~{}\text{vs.}~{}~{}~{}~{}~{}\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\\ ~{}~{}~{}~{}~{}\text{or}&\text{useful}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\\ &\\ \text{M}_{1}(X_{i}),~{}\text{M}_{2}(X_{i})>Y_{i}&\\ ~{}~{}~{}~{}~{}\text{or}&\text{useless}\\ \text{M}_{1}(X_{i}),~{}\text{M}_{2}(X_{i})<Y_{i}&\\ \end{cases}$ (2)
This analysis reveals a potential weakness in benchmarks that rely on anchored comparisons to reference texts. If the reference texts are not carefully calibrated in terms of difficulty, the ranking of LLMs could be determined by only a small portion of the test prompts. Specifically, if reference texts are consistently superior or inferior to most model outputs, a large number of comparisons become uninformative. This issue is mitigated when directly comparing model outputs without a reference text, suggesting that direct comparisons may provide a more robust method for ranking LLMs.
3.3 Iterated Tournaments as a Rough Approximation of Merge Sort
In this section, we argue that iterated tournaments will reach to a good sorting of participants LLMs comparing the similarities in merge sort algorithm and iterated tournament regarding the presumptions of Elo model for LLM capability.
Our approach is predicated on a key assumption: although the prompts in the benchmark dataset are diverse, they can be considered approximately equivalent in their capacity to differentiate LLM performance. This assumption of ‘indistinguishable prompts’ aligns with the Elo model of LLM capability, which posits that prompt-specific variations should not significantly affect the expected outcome likelihood of a single match. [^1] Given the simplicity of the Elo model’s assumptions regarding LLM capability, we posit that ranking LLMs with Elo ratings for a mixture of disparate capabilities may be inappropriate, as a higher Elo rate does not guarantee uniform mastery across multiple capabilities.
Building on this assumption, we propose that conducting tournaments across the benchmark prompts can lead to an accurate ranking of LLMs. We suggest that single-elimination tournaments function analogously to merge sort, with a crucial distinction: they perform only the initial comparison at each step of a merge operation and discard the losing side (i.e. only the winner proceeds to the next rounds). This makes a single tournament perform $n-1$ comparisons in total. This number of comparisons is less than $\mathcal{O}(n\text{log}_{2}n)$ which most of the efficient sorting algorithms requires. However, regarding the assumption of ‘indistinguishable prompt’ we mentioned above, number of tournament matches from the other dataset prompt would help filling the lack of comparion matches. By iterating over the benchmark dataset with randomized initial match-ups, these tournament matches can eventually approximate the complete set of comparisons required for merge sort, given a sufficient number of benchmark prompts ($|X|$).
Our approach provides $\mathcal{O}(|X|$/$\text{log}_{2}n_{\text{model}})$ times more comparisons [^2] Where $\mathcal{O}(|X|$/$\text{log}_{2}n_{\text{model}})=\frac{|X|(n_{\text{model}}-1)}{\mathcal{O}(n_{\text{model}}\text{log}_{2}n_{\text{model}})}$ representing the ratio between matches in iterated tournaments and matches needed for sorting LLMs with an ideal judge. than the minimum required for merge sorting $n_{\text{model}}$ LLMs. In practice, number of matches to figure which model is better from a pair is much larger since the judge model will not be ideal and also performance of LLMs will be stochastic regarding the prompt it was asked to answer. Nevertheless, the large number of prompts ($|X|$), typically ranging from hundreds to thousands, lead us to the worth of empirical validation of the idea, that iterated tournaments will accurately sort LLMs with confidence. This foundational argument underpins our unbiased judge simulation experiment to test the potential of tournament approach (Varco Arena).

### v2

@@ Abstract
Most existing benchmarking approaches for evaluating the output quality of large language models (LLMs) rely on comparing LLM responses to predefined references. Such methods, based on static datasets, quickly become outdated as LLM capabilities and use cases evolve. In this work, we introduce VARCO Arena—a novel, cost-effective, and robust benchmarking approach that leverages a single-elimination tournament structure to minimize the number of required comparisons while eliminating the need for static references or costly human annotations. We validate our approach through two experiments: (i) a simulation study that examines its robustness under various conditions, and (ii) an empirical evaluation using publicly available benchmark prompts. In both experiments, VARCO Arena consistently outperforms current LLM benchmarking practices, achieving stronger correlations with human-established Elo ratings. Our results demonstrate that VARCO Arena not only produces reliable LLM rankings but also provides a scalable, adaptable solution for qualitative evaluation across diverse, customized use cases. We release our demo and code at [URL placeholder].
VARCO Arena: A Tournament Approach to Reference-Free Benchmarking Large Language Models
Seonil Son1, Ju-Min Oh1thanks: Current institution: Samsung AI Research, Heegon Jin1thanks: Current institution: Global AI Platform, Cheolhun Jang1, Jeongbeom Jeong1, Kuntae Kim1 1NCSOFT Correspondence: deftson@ncsoft.com
@@ 1 Introduction
The versatility of Large Language Models (LLMs) stems from their generative capacity to address a wide array of tasks. The multi-faceted capability of LLMs enables flexible applications across numerous user scenarios Ouyang et al. (2022); Köpf et al. (2024); Roziere et al. (2023). As versatility emerges as a core attribute of LLMs, the challenge of accurately gauging their skill becomes increasingly significant. In response to this challenge, numerous benchmarks evaluating LLM capabilities have emerged Hendrycks et al. (2020); Srivastava et al. (2023); Zhong et al. (2024). Many LLM benchmarks employ formats amenable to automated scoring. Examples include benchmarks for arithmetic problems Gao et al. (2022); Patel et al. (2021), multiple-choice questions Lin et al. (2022), and code execution Austin et al. (2021); Chen et al. (2021). While these benchmarks are valuable, they primarily assess problem-solving abilities. The need for LLM benchmarks extends beyond this, as the primary value of LLMs lies in the versatility of their generative behavior. Researchers have primarily relied on pairwise comparisons to evaluate and rank the LLMs based on human preference annotations. A representative example of this approach is Chatbot Arena Chiang et al. (2024), which computes Elo ratings based on a massive number of human votes. While Chatbot Arena benefited from reliable, open-ended prompts generated by a large user base, obtaining such a high volume of annotations from dedicated annotators remains challenging.
@@
Figure 1: VARCO Arena directly compares LLM response pairs in single-elimination tournament rather than comparing references. In terms of deciding whether a certain LLM is better or worse compared to the other one, we suggest direct head-to-head comparison is more intuitive and results in better separability.
No. Comp. ($\downarrow$) Judge Eval. Type
Chatbot Arena unknown human head-to-head
Current Practice $n_{\text{model}}\cdot|X|$ LLM reference -based
VARCO Arena $(n_{\text{model}}-1)\cdot|X|$ LLM head-to-head
Table 1: Comparison between Current Practice and Varco Arena. $|X|$ and $n_{\text{model}}$ represents size of benchmark dataset, and number of candidate LLMs to rank respectively. Human annotators are considered much more costly than LLM judge counterpart.
To address these challenges, several benchmark datasets for generation tasks have been developed and applied with an LLM judge Zheng et al. (2024) to quantify LLM capabilities. Using these benchmarks and their reference responses, LLMs are ranked via pairwise comparisons between their outputs and the reference responses provided by LLM judges. Representative examples include AlpacaEval Li et al. (2023), Arena-Hard-Auto Li et al. (2024), and MTBench Zheng et al. (2023). There are two major advantages to relying on reference responses: (1) The number of comparisons required for ranking scales linearly with the number of LLMs. (2) The reference response establishes a quality standard for evaluating LLM outputs.
However, we argue that using reference responses to mediate comparisons between LLM outputs is suboptimal and that head-to-head comparisons yield more accurate assessments. In this context, we propose VARCO Arena as a novel evaluation framework that directly compares LLM responses while requiring fewer comparisons. Even with fewer comparisons and without relying on reference responses, VARCO Arena achieves a stronger correlation with the human-established Elo rankings from Chatbot Arena.
VARCO Arena is designed to conduct a single-elimination tournament for each prompt across all participating LLM responses and then compute Elo ratings Elo and Sloan (1978) based on the outcomes of these matches. By organizing tournament-style matches for multiple LLMs on each prompt, we obtain relative win rates between all possible model pairs with fewer comparisons than those required by reference-based methods. These win rates are subsequently interpreted as Elo ratings, enabling us to generate a comprehensive ranking of candidate LLMs.
We validate the effectiveness of VARCO Arena against current LLM benchmarking practices through two experiments. First, we conduct a simulation study (Section 4.2) to evaluate the reliability and robustness of our tournament approach under various controlled factors, including the number of participating LLMs, the number of test prompts, and the accuracy of the judge model. Second, we empirically demonstrate that running VARCO Arena with several judge LLMs on public benchmark (Section 4.3) prompts yields a stronger correlation with human-established Elo rankings from Chatbot Arena than does the reference-based approach.
These experiments confirm that our benchmarking approach outperforms current practices in terms of ranking reliability while requiring the lesser number of comparisons (Table 1). Notably, the simulation results—which emulate outcomes under various conditions—underscore that the superiority of our tournament approach is not coincidental but rather reflects the inherent advantages of direct pairwise comparisons over mediated comparisons using reference responses.
In this work, we make the following contributions:
@@
    We introduce VARCO Arena, a novel reference-free benchmarking method for LLMs that leverages a single-elimination tournament structure to perform head-to-head comparisons of LLM responses across test prompts. This approach eliminates the dependency on static references and enhances benchmarking flexibility.
@@
    Through extensive simulation studies and empirical evaluations, we demonstrate that VARCO Arena not only yields more reliable LLM rankings compared to traditional reference-based methods but also achieves this with significantly fewer comparisons, thereby reducing evaluation costs.
  3. 3.
    We release our demo and code to facilitate the further research and reproducibility; the code is available at [URL placeholder]
@@ 2 Preliminaries: Quantifying Generation Ability
Quantifying an LLM’s generation ability is crucial, but it presents several challenges. The outcomes of comparisons between two LLMs are often probabilistic, influenced by various factors such as the provided test prompts and the inherently subjective nature of human preferences. A straightforward approach is to evaluate an LLM’s performance across diverse test prompts, which approximates its real-world performance. Two popular measures for LLM generation ability are the win rate against reference responses and the Elo rating.
2.1 Win Rate in Comparison to the Reference
AlpacaEval Li et al. (2023) and Arena-Hard-Auto Li et al. (2024) are representative benchmark datasets that aim to quantify LLM response quality using reference responses. These benchmarks employ an LLM judge with a standardized prompt for automated evaluation of generation capability. Given the prompts, LLM responses, and reference responses, the judge LLM is tasked with determining whether the reference or the response is preferred. The LLM’s win rate across the test prompts in the benchmark is used as a measure of its generation proficiency.
@@ 2.2 Elo Rating
The Elo rating system, introduced by Elo and Sloan, has since become a popular method for quantifying performance levels in competitive sports and, more recently, for evaluating the generative abilities of LLMs. The primary purpose of the Elo rating system is to represent a participant’s skill level as a single scalar value, enabling the prediction of relative win rates between participants who have never directly competed. Assessing the superiority of one LLM over another in terms of generative capability shares several similarities with determining winners in competitive sports. Varying test prompts may yield different results, similar to how weather or other factors can influence the outcome of a sports match. While a superior team or LLM is not guaranteed to outperform its inferior counterpart in every instance, it tends to succeed more frequently. With Elo ratings, one can expect the relative chance of winning between a pair of players. The expected win rate of the player $i$ against $j$ is computed as follows:
@@ 3 VARCO Arena
We propose VARCO Arena, a reference-free approach for benchmarking large language models (LLMs) via single-elimination tournaments. Instead of comparing model responses against fixed references, VARCO Arena directly compares outputs from different models, determining superiority through head-to-head matchups for each prompt in our benchmark dataset. Repeated tournaments across prompts yield reliable leaderboards that reflect the relative performance of each model.
We begin by motivating our approach over current reference-based evaluations (Section 3.1). Next, we detail how VARCO Arena performs tournaments and results in Elo ratings (Section 3.2 and Algorithm 1). Finally, we discuss the reason why aggregating the results from tournaments promises reliable ranking (Section 3.3).
3.1 Comparing to a Reference is not Always Helpful
Although reference texts are a standard way to evaluate and rank large language models (LLMs), they introduce potential failure modes. Beyond the fact that a single reference might not capture every dimension of correctness, relying solely on a reference can lead to unreliable rankings of LLMs.
Consider an ideal scenario with a judge capable of perfectly distinguishing the quality of any two outputs. If we choose to compare LLM responses directly to rank them using Elo ratings (Equation 1), all head-to-head comparisons are utilized. In contrast, reference-based evaluation for differentiating LLMs can exhibit failure modes, as shown in Equation 2.
$\begin{array}[]{l}\text{M}_{1}(X_{i})\\ \quad\text{vs.}\,\rightarrow\\ \text{M}_{2}(X_{i})\end{array}\hfil\hskip 20.00003pt&\begin{cases}\text{M}_{1}(X_{i})>Y_{i}>\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i})<Y_{i}<\text{M}_{2}(X_{i})&\text{(helpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})>Y_{i}&\text{(unhelpful)}\\ \text{M}_{1}(X_{i}),\ \text{M}_{2}(X_{i})<Y_{i}&\text{(unhelpful)}\end{cases}\end{array}$
When the reference output ($Y_{i}$) for a prompt ($X_{i}$) successfully disambiguates the pair of LLM responses $M_{1}(X_{i})$ and $M_{2}(X_{i})$ (as in the first and second cases), comparison to the reference is effective for benchmarking. Otherwise, these comparisons do not help differentiate LLM performance. Consequently, the reference-based approach provides less information for ranking when multiple responses are either both correct or both incorrect relative to the reference.
3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
Algorithm 1 Tournaments of LLMs over prompts
1:prompts $X=\{x_{1},~{}x_{2},~{}...,~{}x_{i}\}$, LLMs $M=\{m_{1},~{}m_{2},...,~{}m_{j}\}$, outputs $O_{i,j}$ = $m_{j}(x_{i})$
@@ 3.2 Tournaments of LLMs over multiple prompts to Elo Ratings
4:     return $m_{1}$ if $\texttt{IsBetter}(O_{x,1},O_{x,2})$
5:     else $m_{2}$
6:end function
7:function SingleElim($M$, $x$, res)
8:     if $|M|=2$ then
9:         res.append(Match($M$[0], $M$[1], $x$))
10:         return res[-1]
11:     end if
12:     mid $\leftarrow\lfloor|M|/2\rfloor$
13:     left $\leftarrow$ SingleElim($M$[:mid], $x$, res)
14:     right $\leftarrow$ SingleElim($M$[mid:], $x$, res)
15:     return SingleElim(left + right, $x$, res)
16:end function
17:function Tournaments2Ranks($X,M$)
@@
19:     for $x_{i}\in X$ do
20:         SingleElim(Shuffled($M$), $x_{i}$, res)
21:     end for
22:     return ComputeElo(res)
23:end function
Figure 1 and Algorithm 1 illustrate how VARCO Arena benchmarks LLMs via a tournament approach. Here, $|X|$ denotes the number of prompts in the benchmark dataset. Each execution of VARCO Arena runs a tournament among participant LLMs for every prompt in the dataset.
The use of tournament structures for LLM benchmarking offers both benefits and challenges. A major advantage of a single-elimination tournament is efficiency. As shown in Table 1, the number of matches scales linearly with the number of participants and even lower compared to using references. However, single elimination tournament only identifies a champion, leaving the relative ordering of other participants unclear.
To retain tournament’s efficiency while obtaining a fine-grained ranking, we propose aggregating tournament results over multiple prompts with randomized initial match-ups for each prompt. Performing multiple tournaments with random initialization offers several benefits:
@@
    It resolves ties among non-champion participants from previous tournaments.
@@
    It mitigates the impact of unfavorable match-ups in any single tournament.
@@
    Aggregating match results allows for precise win rate estimation via Elo ratings, resulting in a well-aligned overall ranking.
  4. 4.
    More matches are allocated to high-performing participants while ensuring every participant is evaluated at least once per prompt.
In Section 3.3, we further explain how aggregating multiple tournaments could yield an reliable ranking of LLMs. We also provide an analysis of the number of matches each LLM faces, offering a comprehensive view of the method’s efficiency and effectiveness.
3.3 Why Aggregating Multiple Tournaments Yields Reliable Ranks?
Our approach aggregates match outcomes from multiple tournaments to approximate the complete set of pairwise comparisons—akin to the comparisons made in merge sort. In a single-elimination tournament, every participant advances based solely on match outcomes, a process that mirrors the merging steps in merge sort. Notably, a single-elimination tournament executes only the comparisons strictly necessary for determining the winner, omitting many comparisons that merge sort would perform further.
We posit that the missing pairwise match-ups in any one tournament can be recovered by aggregating tournaments conducted over different prompts. This hypothesis relies on the assumption—central to our use of the Elo model—that match outcomes are independent of the prompt. Consequently, matches across different prompts are considered equivalent.
Considering only the initial comparisons, which are randomly sampled, the aggregate number of comparisons is $|X|\cdot n_{\text{model}}/2$. Since $n_{\text{model}}$ is typically on the order of tens and $|X|$ comprises at least hundreds, this number exceeds the total possible match-ups, $\binom{n_{\text{model}}}{2}$ [^1] Note that merge sort produces a set of match-ups with no duplicates.. Furthermore, as shown in Table 1, the remaining matches—totaling $|X|\cdot(n_{\text{model}}-1)$—either contribute additional merge sort structures or enhance the accuracy of estimating relative win rates among LLM participants.
Moreover, considering that each unique pair meets in at least $|X|/(2(n_{\text{model}}-1))$ matches [^2] A model participates in between $|X|$ and $\lceil\log_{2}n_{\text{model}}\rceil$ matches per tournament. Dividing this by the number of possible match-ups per model, $2(n_{\text{model}}-1)$, yields the expression above. across the benchmark, this frequency is sufficient for a fair estimation of their relative win rates.
In summary, aggregating tournaments not only reconstructs the full set of pairwise comparisons for a merge sort but also ensures that each pair of models faces one another often enough to yield accurate win rate estimations, leading to reliable Elo ratings. Based on these considerations, we propose that conducting tournaments over a benchmark prompts will yield a reliable ranking of LLMs.